OpenStudy (anonymous):

Help solve : (3x-2)/(x+4) = -5

OpenStudy (anonymous):

cross multiply

OpenStudy (anonymous):

how so ?

OpenStudy (anonymous):

(3x - 2) = -5 (x+4) 3x - 2 = -5x - 20 8x = -18 x = -18/8 = -9/4